# Boundary phase transitions of two-dimensional quantum critical XXZ model

### v1

@@ Abstract
The boundary critical behavior of the two-dimensional (2D) quantum antiferromagnetic (AF) XXZ model coupled with either a dangling spin-$1/2$ XXZ chain or a dangling two-leg ladder on the boundary is studied with the bosonization and renormalization group analysis. A rich boundary phase diagram is obtained in each case. In the dangling chain case, the boundary either develops a long-range AF order in the easy-plane or the easy-axis direction with extraordinary critical behavior, or has a valence bond solid order with ordinary boundary critical behavior. In the case of a dangling two-leg ladder, besides the possible easy-plane or easy-axis AF ordered phases on the boundary with extraordinary critical behavior, the boundary can form a singlet phase with ordinary critical behavior without breaking any symmetry.
1 Introduction
@@
The universality class of boundary critical behavior can be changed by adjusting the form and strength of boundary interactions, which results into a rich boundary phase diagram. In the classical O($N$) model of the Wilson-Fisher universality, this can be achieved by tuning the interaction strength on the boundary. If the boundary interaction strength is weaker than or similar to the bulk, it inherits long-range order only below the bulk transition temperature $T_{c,b}$, and the boundary critical behavior at $T_{c,b}$ is dubbed the ordinary transition. On the other hand, strong interactions on the boundary can induce long-range order at higher temperature than $T_{c,b}$ if the boundary itself is above the lower critical dimension, then the singular behavior of boundary physical quantities at $T_{c,b}$ is dubbed the extraordinary transition. The multicritical point separating the ordinary and the extraordinary phases is called the special transition. These boundary universality classes have been extensively studied with analytical and numerical methods for the classical O($N$) models [1, 2, 12, 13].
@@
In 2D quantum spin models with O(3) symmetry, the quantum phase transition from a topological phase protected by the spin rotation and lattice translation symmetry to the long-range antiferromagnetic (AF) order can be realized in the spin-$1/2$ Heisenberg model on decorated square lattice [8] and the spin-$1$ coupled Haldane chains [15]. While the QCPs belong to the (2+1)D O(3) Wilson-Fisher universality class in the bulk, the boundary critical behavior in both models is distinct from the ordinary universality. Moreover, various classes of nonordinary boundary critical behavior have been found in numerical simulations of other quantum spin models with dangling spin chains or ladders [16, 17, 18, 19, 20, 21]. This inspires further theoretical analysis of the boundary critical behavior of (2+1)D O($N$) universality.
In Ref. [22], the boundary critical behavior of 3D O($N$) model coupled with a 2D system on the boundary is studied with RG analysis. The boundary is described by a 2D O($N$) nonlinear $\sigma$ model. Starting from the “normal” boundary state polarized by an infinitesimal external field, it is shown that its coupling to the bulk suppresses the gapless order parameter fluctuations on the boundary, thus stabilizes the weak-coupling fixed point. Moreover, the consequent slow RG flow induces a logarithmic decay of spin correlations instead of a long-range order on the boundary, which is dubbed the extraordinary-log universality class. This peculiar behavior is observed in numerical simulations of 3D classical spin models with O(3) [23, 24] or O(2) symmetry [25, 26] and also in the 3D AF Potts model with an emergent O(2) symmetry at the critical point [27, 28].
In Ref. [29], the boundary critical behavior of a (2+1)D O(3) QCP coupled with a spin-$1/2$ AF Heisenberg chain is studied. The dangling spin chain is treated as a quantum critical state with the non-Abelian bosonization. Its coupling to the bulk state leads to two possible phases on the boundary: one has the valence bond solid (VBS) order spontaneously breaking the translation symmetry, while the other is argued to have long-range AF order, and there is a direct phase transition between them. The nonordinary boundary critical behavior found in numerical simulations is argued to correspond to the second phase with a weak AF order, which is confirmed in later numerical simulations [30]. Moreover, weak AF order is also observed on the boundary of a 2D quantum critical Heisenberg model with a dangling two-leg ladder [20], and a quantum XXZ model with a dangling chain [21].
@@
In this work, we theoretically study the boundary critical behavior of the 2D quantum critical XXZ model with either a dangling spin-$1/2$ chain or a dangling two-leg ladder, which are illustrated in Fig. 1. In each case, we treat the dangling spin chain or the ladder with the Abelian bosonization method and take its coupling to the bulk as perturbations in the RG analysis. We find a rich boundary phase diagram in each case. The results depend on whether the system has an easy-plane or an easy-axis anisotropy, and we particularly focus on the easy-plane case, which reduces to the O(3) symmetric case if the anisotropy is turned off.
In the dangling spin chain case with easy-plane anisotropy, we find a boundary phase diagram similar to the O(3) symmetric case studied in Ref. [29]. The boundary has either a VBS order with ordinary critical behavior, or a long-range AF order with extraordinary critical behavior. In the latter case, the long-range AF order is shown based on the solution of the Callan-Symanzik equation of the spin correlation function. In the dangling ladder case, a spin-singlet phase with ordinary critical behavior without breaking any symmetry emerges besides the AF order phases on the boundary. Therefore, the RG analysis is consistent with the boundary phases observed in numerical simulations of the quantum XXZ and O(3) models.
The rest of the paper is organized as follows. The boundary critical behavior of the 2D quantum critical XXZ model with a dangling spin chain and with a dangling two-leg ladder is studied in Sec. 2 and 3, respectively. In each section, we first review the bosonization and the quantum phase diagram of the spin chain or the two-leg ladder for completeness, then we consider the coupling with the bulk and the boundary phase diagram of the coupled system. We consider the effect of weak velocity difference between the bulk and the boundary modes in Sec. 4, and conclude in Sec. 5.
2 Boundary critical behavior with a dangling spin chain
In this section, we study the 2D quantum XXZ model with a dangling spin-$1/2$ chain on the boundary, which is illustrated in Fig. 1 (a). We first review the bosonization and the phase diagram of the XXZ chain for completeness, then consider its coupling to the bulk critical state and study the phase diagram of the coupled system.
2.1 Spin-$1/2$ XXZ chain
@@
Here, the case with $\Delta=0$ corresponds to free fermions with Fermi momentum $k_{\mathrm{F}}=\pi/(2a_{0})$, in which $a_{0}$ is the lattice constant.
@@
In the bosonization dictionary [31], the fermion field operators are expressed by
@@
Linearizing the free fermion dispersion relation near the Fermi points, and using the bosonization dictionary [31], the action of the spin-$1/2$ XXZ chain is cast into the following sine-Gordon model (see Appendix A for calculation details),
@@
2.2 Coupling to the bulk
@@
in which $\epsilon_{a}$ is the boundary anomalous dimension of the ordinary universality class, which can be calculated perturbatively with field-theoretic methods [2]. Here, it is implicitly assumed that the velocity of the bulk critical mode is the same as the dangling spin chain. We will show that a weak velocity difference is irrelevant in the RG flow in Sec. 4.1.
@@
which reduces to the RG equations of the XXZ spin chain in Eqs. (19)–(20) if the couplings $g_{xx}$ and $g_{z}$ are turned off. Solving the coupled RG equations completely in the high-dimensional parameter space is in general quite difficult, thus we focus on the influence of weak couplings $g_{xx}$ and $g_{z}$ on the phases of the spin chain.
Starting from the VBS order phase of the spin chain, in which $\lambda\rightarrow-\infty$ and $K\rightarrow 0$ in the RG flow, we find both coupling strengths $g_{xx}$ and $g_{z}$ are renormalized to zero, thus the dangling spin chain and the critical bulk mode effectively decouple from each other. Therefore, the the VBS order and the ordinary boundary critical behavior coexist on the boundary.
Starting from the Ising-AF order phase of the spin chain, where $\lambda\rightarrow+\infty$ and $K\rightarrow 0$ in the RG flow, we find that $g_{xx}$ is renormalized to zero while $g_{z}$ diverges in the RG flow, and it turns out to accelerate the flow of $\lambda$ and $K$ towards infinity and zero, respectively. Therefore, the Ising-AF order on the boundary is stabilized, and induces an extraordinary boundary spin correlations in the $z$ direction. In contrast, the boundary spin correlations in the $xy$ direction retains its ordinary critical behavior, because the coupling $g_{xx}$ is renormalized to zero.
@@
with the initial conditions $K(l=0)=K_{0}$ and $g_{xx}(l=0)=g_{0}$. The simplified RG equations have the same form as those of the sine-Gordon model and can be exactly solved. In order to focus on the asymptotic behavior of the solution in the $K\rightarrow+\infty$ regime, we drop the subleading $1/(4K)$ term in Eq. (28), and find the asymptotic solution,
$K(l)\simeq\frac{\pi^{2}g_{0}^{2}}{1+2\epsilon_{xx}}\big{(}e^{(1+2\epsilon_{xx})l}-1\big{)}+K_{0},\quad g_{xx}(l)=g_{0}e^{(1/2+\epsilon_{xx})l}.$ (29)
The two-point correlation function of a given field operator in the momentum space $G^{(2)}(p;g_{xx})$ satisfies the Callan-Symanzik equation,
@@
From the correlation functions of the boson fields $\phi$ and $\theta$ in the free theory, Eqs. (12) and (13), we find the factors $Z_{\phi}=K_{0}/K$ and $Z_{\theta}=K/K_{0}$, thus we find $\gamma_{\phi}=-\pi^{2}g_{xx}^{2}/(2K)$, and $\gamma_{\theta}=\pi^{2}g_{xx}^{2}/(2K)$. The two-point correlation functions are obtained by integrating the Callan-Symanzik equation along the RG flow trajectory detailed in Appendix C,
@@
$A=\frac{2^{-4\epsilon_{xx}-3}\pi}{\sin(2\epsilon_{xx}\pi)\Gamma(2\epsilon_{xx}+2)^{2}},\quad B=-\frac{2^{4\epsilon_{xx}+1}\sin(2\epsilon_{xx}\pi)\Gamma(2\epsilon_{xx}+1)^{2}}{\pi}$ (39)
@@
$K=\frac{1}{2}-\epsilon_{xx},\,\lambda=-\frac{\epsilon_{xx}+\epsilon_{z}}{\pi},\,g_{xx}=\frac{\sqrt{2(\epsilon_{xx}+\epsilon_{z})(3\epsilon_{xx}+\epsilon_{z})}}{\pi},\,g_{z}=\frac{\sqrt{8\epsilon_{xx}(\epsilon_{xx}+\epsilon_{z})}}{\pi}$ (42)
up to the first order in $\epsilon_{xx}$ and $\epsilon_{z}$, which controls the transition from the VBS order to the AF order on the boundary. In particular, if $\epsilon_{xx}=\epsilon_{z}$, the fixed point satisfies $g_{xx}=g_{z}$ thus has the O(3) symmetry, and corresponds to the one found in Ref. [29]. Linearizing the RG equations near the fixed point and finding the scaling dimensions of the eigenoperators under an infinitesimal scale transformation, we find two scaling dimensions are positive. This indicates that the new fixed point has two relevant directions, thus the boundary transition from the AF order to the VBS order in the XXZ model with a dangling spin chain is in general first-order. However, in the O(3) symmetric case, only one relevant boundary operator is allowed by the O(3) symmetry, thus the boundary transition between the AF order and the VBS order is continuous, which has been proposed in Ref. [29] and observed in recent numerical simulations [30].
3 Boundary critical behavior with a dangling two-leg ladder
In this section, we study the boundary critical behavior of the 2D quantum XXZ model with a dangling two-leg ladder, which is illustrated in Fig. 1 (b). The same strategy as in the dangling spin chain case is adopted. We first depict the phase diagram of the two-leg ladder, then examine the stability of each phase against a weak coupling to the critical bulk states.
3.1 Two-leg ladder

### v3

@@ Abstract
The boundary critical behavior of the two-dimensional (2D) quantum antiferromagnetic (AF) XXZ model coupled with either a dangling spin-$1/2$ XXZ chain or a dangling two-leg ladder on the boundary is studied with the bosonization and renormalization group analysis. A rich boundary phase diagram is obtained in each case. In the dangling chain case, the boundary either develops a long-range AF order in the easy-plane or the easy-axis direction with extraordinary critical behavior, or has a valence bond solid order with ordinary boundary critical behavior. In the case of a dangling two-leg ladder, besides the possible easy-plane or easy-axis AF ordered phases on the boundary with extraordinary critical behavior, the boundary can form a singlet phase with ordinary critical behavior without breaking any symmetry. These results are consistent with recent numerical simulations on the boundary critical behavior of 2D quantum spin models.
I Introduction
@@
The universality class of boundary critical behavior can be changed by adjusting the form and strength of boundary interactions, which results into a rich boundary phase diagram [1]. In the classical O($N$) model of the Wilson-Fisher universality, this can be achieved by tuning the interaction strength on the boundary. If the boundary interaction strength is weaker than or similar to the bulk, it inherits long-range order only below the bulk transition temperature $T_{c,b}$, and the boundary critical behavior at $T_{c,b}$ is dubbed the ordinary transition. On the other hand, strong interactions on the boundary can induce long-range order at higher temperature than $T_{c,b}$ if the boundary itself is above the lower critical dimension, then the singular behavior of boundary physical quantities at $T_{c,b}$ is dubbed the extraordinary transition. The multicritical point separating the ordinary and the extraordinary phases is called the special transition. These boundary universality classes have been extensively studied with analytical and numerical methods for the classical O($N$) models [1, 2, 12, 13].
@@
In 2D quantum spin models with O(3) symmetry, the quantum phase transition from a topological phase protected by the spin rotation and lattice translation symmetry to the long-range antiferromagnetic (AF) order can be realized in the spin-$1/2$ Heisenberg model on decorated square lattice [8] and the spin-$1$ coupled Haldane chains [15]. While the QCPs belong to the (2+1)D O(3) Wilson-Fisher universality class in the bulk, the boundary critical behavior in both models is distinct from the ordinary universality. Moreover, various classes of nonordinary boundary critical behavior have been found in numerical simulations of other quantum spin models with dangling spin chains [16, 17, 18, 19]. This inspires further theoretical analysis of the boundary critical behavior of (2+1)D O($N$) universality.
In Ref. [20], the boundary critical behavior of 3D O($N$) model coupled with a 2D system on the boundary was studied with RG analysis. The boundary is described by a 2D O($N$) nonlinear $\sigma$ model. Starting from the “normal” boundary state polarized by an infinitesimal external field, it is shown that its coupling to the bulk suppresses the gapless order parameter fluctuations on the boundary, thus stabilizes the weak-coupling fixed point. Moreover, the consequent slow RG flow induces a logarithmic decay of spin correlations instead of a long-range order on the boundary, which is dubbed the extraordinary-log universality class. This peculiar behavior has been observed in numerical simulations of 3D classical spin models with O(3) [21, 22] or O(2) symmetry [23, 24] and also in the 3D AF Potts model with an emergent O(2) symmetry at the critical point [25, 26].
In Ref. [27], the boundary critical behavior of a (2+1)D O(3) QCP coupled with a spin-$1/2$ AF Heisenberg chain was studied. The dangling spin chain is treated as a quantum critical state with the non-Abelian bosonization. Its coupling to the bulk state leads to two possible phases on the boundary: one has the valence bond solid (VBS) order that spontaneously breaks the translation symmetry, while the other is argued to have long-range AF order, and there is a direct phase transition between them. The nonordinary boundary critical behavior found in numerical simulations was argued to correspond to the second phase with a weak AF order, which has been confirmed by latest numerical simulations [28]. Moreover, weak AF order was also observed on the boundary of the easy-plane XXZ model with a dangling spin chain [see Fig. 1 (a)] [29], indicating that the extraordinary boundary critical behavior is not restricted to the isotropic case.
Recently, the boundary critical behavior of spin-$1/2$ Heisenberg and XXZ models with a dangling two-leg ladder [see Fig. 1 (b)] was revisited [30, 29]. At the bulk QCP, the dangling ladder weakly coupled to the bulk exhibits the ordinary critical behavior, which is in the same universality class of the 3D classical O($N$) ($N=2,3$) model [16, 30, 29]. However, increasing the boundary interaction strength induces a phase transition to a weakly AF ordered state on the boundary [30]. This indicates a nontrivial interplay of the quasi-1D dangling ladder and the critical bulk state, and has not been expected before. This motivates our theoretical analysis in this work.
@@
In this work, we theoretically study the boundary critical behavior of the 2D quantum critical XXZ model coupled to a dangling spin-$1/2$ chain or a dangling two-leg ladder, which are illustrated in Fig. 1. In each case, we treat the dangling spin chain or the ladder with the Abelian bosonization method and take its coupling to the bulk as a perturbation in the RG analysis. We find a rich boundary phase diagram in each case. The results depend on whether the system has an easy-plane or an easy-axis anisotropy, and we particularly focus on the easy-plane case, which reduces to the O(3) symmetric case if the anisotropy is turned off.
In the dangling spin chain case with easy-plane anisotropy, we find a boundary phase diagram similar to the O(3) symmetric case studied in Ref. [27]. The boundary has either a VBS order with ordinary critical behavior, or a long-range AF order with extraordinary critical behavior. In the latter case, the long-range AF order is shown based on the solution of the Callan-Symanzik equation of the spin correlation function. In the dangling ladder case, the boundary can be either in a spin-singlet phase with ordinary critical behavior without breaking any symmetry, or in an AF ordered phase with extraordinary critical behavior. Therefore, the RG analysis is consistent with the boundary phases observed in numerical simulations of the quantum XXZ and O(3) models.
The rest of the paper is organized as follows. The boundary critical behavior of the 2D quantum critical XXZ model with a dangling spin chain and with a dangling two-leg ladder is studied in Sec. II and III, respectively. In each section, we first review the bosonization and the quantum phase diagram of the spin chain or the two-leg ladder, then we consider the coupling with the bulk and derive the boundary phase diagram of the coupled system from the coupled RG equations. We consider the effect of weak velocity difference between the bulk and the boundary modes in Sec. IV, and conclude in Sec. V.
II Boundary critical behavior with a dangling spin chain
In this section, we study the 2D quantum XXZ model with a dangling spin-$1/2$ chain on the boundary, which is illustrated in Fig. 1 (a). We first review the bosonization and the phase diagram of the XXZ chain [31, 32] for completeness, then consider its coupling to the bulk critical state and study the phase diagram of the coupled system.
II.1 Spin-$1/2$ XXZ chain
@@
Here, $\Delta=0$ corresponds to free fermions with the Fermi momentum $k_{\mathrm{F}}=\pi/(2a_{0})$, in which $a_{0}$ is the lattice constant.
@@
In the bosonization dictionary [31, 32], the fermion field operators are expressed by
@@
Linearizing the free fermion dispersion relation near the Fermi points, and using the bosonization dictionary [31, 32], the action of the spin-$1/2$ XXZ chain is cast into the following sine-Gordon model (see Appendix A for calculation details),
@@
II.2 Coupling to the bulk
Table 1: Summary of possible boundary phases of the 2D quantum critical AF XXZ model weakly coupled with a dangling spin-$1/2$ XXZ chain on the boundary.
easy-plane AF spin rotation O(2) ordinary extraordinary
VBS lattice translation ordinary ordinary
@@ II.2 Coupling to the bulk
$C_{ab}(z,\bar{z};w,\bar{w})=\langle m_{a}(z,\bar{z})m_{b}(w,\bar{w})\rangle=\frac{\delta_{a,b}}{|z-w|^{3-2\epsilon_{a}}},$ (22)
in which $\epsilon_{a}>0$ ($1\leq a\leq 3$) are the boundary anomalous dimensions of the ordinary universality class, which can be calculated with field-theoretic method [2] and numerical simulations [12]. If the bulk state undergoes a (2+1)D XY [O(2)] transition, the inplane AF mode is critical with an anomalous dimension $\epsilon_{xx}=0.281(2)$ [12]. If the bulk undergoes an Ising transition, the anomalous dimension of the easy-axis critical mode is $\epsilon_{z}=0.2374(15)$ [12]. In the isotropic case with an O(3) transition, the anomalous dimensions $\epsilon_{xx}=\epsilon_{z}=0.313(2)$ [12]. Here, it is implicitly assumed that the velocity of the bulk critical mode is the same as the dangling spin chain. We will show that a weak velocity difference is irrelevant in the RG flow in Sec. IV.1.
@@
which reduces to the RG equations of the XXZ spin chain in Eqs. (19)–(20) if the couplings $g_{xx}$ and $g_{z}$ are turned off. Solving the coupled RG equations completely in the high-dimensional parameter space is in general quite difficult, thus we focus on the influence of weak couplings $g_{xx}$ and $g_{z}$ on each phase of the spin chain below. The results are summarized in Table 1.
Starting from the VBS order phase of the spin chain, in which $\lambda\rightarrow-\infty$ and $K\rightarrow 0$ in the RG flow, we find both coupling strengths $g_{xx}$ and $g_{z}$ are renormalized to zero, thus the dangling spin chain and the critical bulk mode effectively decouple from each other. Therefore, the VBS order and the ordinary boundary critical behavior coexist on the boundary.
Starting from the Ising-AF order phase of the spin chain, where $\lambda\rightarrow+\infty$ and $K\rightarrow 0$ in the RG flow, we find $g_{xx}$ is renormalized to zero while $g_{z}$ diverges in the RG flow, and it turns out to accelerate the flow of $\lambda$ and $K$ towards infinity and zero, respectively. Therefore, the Ising-AF order on the boundary is stabilized, and induces an extraordinary boundary spin correlations in the $z$ direction. In contrast, the boundary spin correlations in the $xy$ direction retains its ordinary critical behavior, because the coupling $g_{xx}$ is renormalized to zero.
@@
with the initial conditions $K(l=0)=K_{0}$ and $g_{xx}(l=0)=g_{0}$. The simplified RG equations have the same form as those of the sine-Gordon model and can be exactly solved. In order to focus on the asymptotic behavior of the solution in the $K\rightarrow+\infty$ regime, we drop the subleading $1/(4K)$ term in Eq. (28), and obtain the asymptotic solution,
$\begin{gathered}K(l)\simeq\frac{\pi^{2}g_{0}^{2}}{1+2\epsilon_{xx}}\big{(}e^{(1+2\epsilon_{xx})l}-1\big{)}+K_{0},\\ g_{xx}(l)\simeq g_{0}e^{(1/2+\epsilon_{xx})l}.\end{gathered}$ (29)
The two-point correlation function in the momentum space $G^{(2)}(p;g_{xx})$ of a given field operator satisfies the Callan-Symanzik equation,
@@
From the correlation functions of the boson fields $\phi$ and $\theta$ in the free theory, Eqs. (12) and (13), we find the factors $Z_{\phi}=K_{0}/K$ and $Z_{\theta}=K/K_{0}$, thus we find $\gamma_{\phi}=-\pi^{2}g_{xx}^{2}/(2K)$, and $\gamma_{\theta}=\pi^{2}g_{xx}^{2}/(2K)$. The two-point correlation functions are obtained by integrating the Callan-Symanzik equation along the RG flow trajectory described in detail in Appendix C,
@@
$A=\frac{2^{-4\epsilon_{xx}-3}\pi}{\sin(2\epsilon_{xx}\pi)\Gamma(2\epsilon_{xx}+2)^{2}},$ (39)
$B=-\frac{2^{4\epsilon_{xx}+1}\sin(2\epsilon_{xx}\pi)\Gamma(2\epsilon_{xx}+1)^{2}}{\pi}$ (40)
@@
$\begin{gathered}K=\frac{1}{2}-\epsilon_{xx},\quad\lambda=-\frac{\epsilon_{xx}+\epsilon_{z}}{\pi},\\ g_{xx}=\frac{\sqrt{2(\epsilon_{xx}+\epsilon_{z})(3\epsilon_{xx}+\epsilon_{z})}}{\pi},\quad g_{z}=\frac{\sqrt{8\epsilon_{xx}(\epsilon_{xx}+\epsilon_{z})}}{\pi}\end{gathered}$ (43)
up to the first order in $\epsilon_{xx}$ and $\epsilon_{z}$, which controls the transition from the VBS order to the AF order on the boundary. In particular, in the isotropic case with $\epsilon_{xx}=\epsilon_{z}$, the fixed point satisfies $g_{xx}=g_{z}$ thus exhibits the O(3) symmetry, which corresponds to the fixed point found in Ref. [27]. Linearizing the RG equations near the fixed point and finding the scaling dimensions of the eigenoperators under an infinitesimal scale transformation, we find two scaling dimensions are positive. This indicates that the new fixed point has two relevant directions, thus the boundary transition from the AF order to the VBS order in the XXZ model with a dangling spin chain is in general first-order. However, in the O(3) symmetric case, only one relevant boundary operator is allowed by the O(3) symmetry, thus the boundary transition between the AF order and the VBS order is continuous, which has been proposed in Ref. [27] and observed in recent numerical simulations [28].
III Boundary critical behavior with a dangling two-leg ladder
In this section, we study the boundary critical behavior of the 2D quantum XXZ model with a dangling two-leg ladder, which is illustrated in Fig. 1 (b). The same strategy as in the dangling spin chain case is adopted. We first review the bosonization of the two-leg ladder and depict its phase diagram for completeness [35, 36, 37, 38, 39, 40, 41, 42], then examine the stability of each phase against a weak coupling to the critical bulk state. We also numerically solve the coupled RG equations with given bare coupling constants as the initial conditions to verify the above perturbative analysis and to explore the nonperturbative regime with relatively strong coupling between the ladder and the bulk state.
III.1 Two-leg ladder